A rectangle has the dimensions 0.7x and 5-3x.What value of x gives the maximum area and what is the maximum area??
Answer by edjones(8007)

The max area of a rectangle is a square.
A square has equal sides.
Hence:
.7x=5-3x
3.7x=5
x=5/3.7 the value of x that gives the maximum area
.
.7(5/3.7)*(5-3(5/3.7))
=(3.5/3.7)*(5-(15/3.7))
=(3.5/3.7)*((18.5-15)/3.7)
=(3.5/3.7)(3.5/3.7) proving both sides are equal.
=(3.5/3.7)^2
=1225/1369 or.8948... units^2
.
Ed